# Quantifying Uncertainty: All We Need is the Bootstrap?

### v2

@@ 3.1 When methods fail to produce a CI
There are only a few cases where a method fails to produce a CI. All bootstrap methods fail to produce a CI for Pearson correlation for $n=4$, due to division by zero variance. It can also happen for $n=8$, but rarely. BC and BCa do not produce a CI for the $5^{th}$ percentile for $n\leq 8$. For the $95^{th}$ percentile, B-t does not produce an interval in most cases when $n=4$ for the Laplace distribution. m-j is unable to produce CIs for small sample sizes and extreme percentiles ($5^{th}$ percentile for $n\leq 16$ and the $95^{th}$ percentile for $n\leq 8$). Method q-nonpar fails to produce the $95^{th}$ percentile for $n=4$ and $\alpha\leq 0.75$, for $n\in\{8,16\}$ and $\alpha\leq 0.25$, for $n=32$ and $\alpha\leq 0.05$ and for $n=64$ and $alpha=0.025$. When predicting CIs for the median, it fails at $n=4$ and $\alpha\leq 0.05$. And, although wilcoxon returns CIs for asymmetric distributions, they are not useful CIs for location.
@@ 3.2 Coverage of bootstrap methods
Table 3.2 shows a comparison of bootstrap methods in mean KL. As expected, coverage improves with sample size and the two extreme percentiles and standard deviation are the most difficult functionals. For Pearson correlation, mean, and median, DB is best. For the percentiles, B-n is best. And for standard deviation, B-t is best. B-n and B-t perform the best overall. Because coverage gets wors with smaller sample size, the overall results are biased towards methods that perform well on small $n$. DB and BCa, which we would expect to perform best, perform relatively poorly for small $n$, but are best and second best for sample sizes $n\geq 64$.
@@
Mean KL coverage performance of bootstrap methods for one-sided CIs. The all column is across all combinations, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std B-n $\underline{0.082}$ $0.322$ $\underline{0.180}$ $\underline{0.091}$ $\underline{0.038}$ $0.021$ $0.014$ $0.010$ $0.006$ $0.030$ $0.004$ $\underline{0.024}$ $\underline{0.145}$ $0.208$ B-t $0.088$ $\underline{0.199}$ $0.280$ $0.112$ $0.051$ $0.024$ $0.020$ $0.014$ $0.008$ $0.015$ $0.012$ $0.078$ $0.264$ $\underline{0.084}$ $B$ $0.117$ $0.253$ $0.194$ $0.227$ $0.106$ $0.047$ $0.034$ $0.023$ $0.017$ $0.034$ $0.045$ $0.159$ $0.234$ $0.145$ SB $0.124$ $0.459$ $0.278$ $0.160$ $0.063$ $0.028$ $0.018$ $0.011$ $0.002$ $0.026$ $0.003$ $0.064$ $0.230$ $0.308$ DB $0.140$ $0.544$ $0.348$ $0.200$ $0.060$ $\underline{0.008}$ $\underline{0.005}$ $\underline{0.003}$ $\underline{0.000}$ $\underline{0.010}$ $\underline{0.002}$ $0.125$ $0.401$ $0.188$ BCa $0.159$ $0.600$ $0.335$ $0.242$ $0.096$ $0.018$ $0.010$ $0.005$ $0.001$ $0.023$ $0.017$ $0.201$ $0.379$ $0.218$ BC $0.162$ $0.607$ $0.350$ $0.234$ $0.096$ $0.022$ $0.013$ $0.008$ $0.001$ $0.026$ $0.017$ $0.169$ $0.387$ $0.246$ PB $0.165$ $0.620$ $0.378$ $0.229$ $0.076$ $0.029$ $0.019$ $0.012$ $0.002$ $0.028$ $0.003$ $0.109$ $0.377$ $0.330$
@@
Threshold-based coverage performance of bootstrapping on one-sided CIs. The table shows the percentage of experiments where a method does not meet the stringent criterion $\text{KL}(0.945,0.95)$. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std DB $\underline{0.64}$ $0.96$ $\underline{0.85}$ $\underline{0.82}$ $\underline{0.73}$ $\underline{0.59}$ $\underline{0.48}$ $\underline{0.20}$ $\underline{0.30}$ $\underline{0.52}$ $0.38$ $0.87$ $\underline{0.78}$ $0.75$ BCa $0.78$ $0.99$ $0.91$ $0.89$ $0.81$ $0.72$ $0.64$ $0.59$ $0.43$ $0.60$ $0.85$ $0.97$ $0.88$ $\underline{0.69}$ B-n $0.78$ $0.97$ $0.93$ $0.89$ $0.82$ $0.71$ $0.67$ $0.59$ $0.67$ $0.76$ $0.65$ $\underline{0.71}$ $0.80$ $1.00$ SB $0.79$ $0.96$ $0.94$ $0.83$ $0.78$ $0.74$ $0.71$ $0.66$ $0.67$ $0.75$ $\underline{0.33}$ $0.98$ $0.97$ $1.00$ PB $0.82$ $0.94$ $0.97$ $0.94$ $0.84$ $0.73$ $0.72$ $0.68$ $0.70$ $0.76$ $0.46$ $0.97$ $0.98$ $1.00$ BC $0.84$ $0.99$ $0.94$ $0.94$ $0.91$ $0.83$ $0.75$ $0.61$ $0.43$ $0.67$ $0.85$ $0.86$ $1.00$ $0.89$ B-t $0.85$ $\underline{0.89}$ $0.88$ $0.87$ $0.87$ $0.84$ $0.84$ $0.81$ $0.80$ $0.64$ $0.91$ $0.92$ $0.95$ $0.88$ $B$ $0.88$ $0.93$ $0.96$ $0.92$ $0.89$ $0.83$ $0.84$ $0.80$ $0.77$ $0.72$ $0.93$ $0.94$ $0.97$ $0.86$
@@
Threshold-based coverage performance of bootstrapping on two-sided CIs. The table shows the percentage of experiments where a method does not meet the liberal criterion $25\times\text{KL}(0.945,0.95)$. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std DB $\underline{0.33}$ $0.81$ $\underline{0.50}$ $\underline{0.58}$ $\underline{0.52}$ $\underline{0.10}$ $\underline{0.03}$ $\underline{0.03}$ $\underline{0.00}$ $0.25$ $\underline{0.07}$ $0.43$ $\underline{0.40}$ $\underline{0.57}$ SB $0.45$ $1.00$ $0.67$ $0.73$ $0.63$ $0.24$ $0.13$ $0.08$ $0.10$ $0.46$ $0.14$ $0.43$ $0.42$ $0.82$ B-n $0.46$ $0.97$ $0.81$ $0.76$ $0.56$ $0.29$ $0.13$ $0.08$ $0.40$ $0.48$ $0.23$ $\underline{0.35}$ $0.43$ $0.80$ BCa $0.48$ $1.00$ $0.89$ $0.69$ $0.63$ $0.35$ $0.11$ $0.06$ $\underline{0.00}$ $0.46$ $0.25$ $0.62$ $\underline{0.40}$ $0.74$ PB $0.48$ $1.00$ $0.78$ $0.77$ $0.63$ $0.32$ $0.13$ $0.08$ $0.20$ $0.49$ $0.19$ $0.47$ $0.47$ $0.82$ BC $0.49$ $1.00$ $0.92$ $0.74$ $0.61$ $0.39$ $0.10$ $0.06$ $\underline{0.00}$ $0.48$ $0.25$ $0.57$ $0.45$ $0.77$ B-t $0.67$ $\underline{0.50}$ $0.72$ $0.81$ $0.76$ $0.66$ $0.61$ $0.60$ $0.40$ $\underline{0.17}$ $0.76$ $1.00$ $1.00$ $0.65$ $B$ $0.82$ $1.00$ $1.00$ $0.94$ $0.85$ $0.79$ $0.68$ $0.63$ $0.60$ $0.49$ $0.95$ $1.00$ $1.00$ $0.79$
@@
From the six confidence levels in our experiments we derive results for the two most common two-sided CIs (95%, 90%). Table 3.2 shows how often the methods fail to meet the liberal criterion for these two-sided intervals. These and results for other thresholds and KL are similar for one-sided and two-sided intervals. Most bootstrap methods meet the liberal criterion in almost all cases when $n\geq 128$ ($n\geq 64$ for DB), but even at $n=256$ there are still two cases where even DB does not (standard deviation for log-normal). Excluding these, DB meets the very liberal criterion for all experiments for $n\geq 64$.
@@ 3.3 Coverage comparison with baseline methods
Note that all the results in this section are for one-sided CIs. Results for two-sided CIs are similar.
We start with the premise that if we are limited to a single method, DB is the best choice. Now we take a closer look at where DB is clearly outperformed by another method. The criterion we use is that method A outperforms method B if B does not meet the stringent criterion and A is at least an order of magnitude better than method B. If a method is not able to produce CIs, it is outperformed by any method that can.
@@
Cases where DB and baseline outperform each other. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6. n functional baseline $\gg$ DB DB $\gg$ baseline 4 corr fisher (6) 4 mean t-test (8); wilcoxon (1) DB (4) 4 median m-j (8); q-par (3) 4 Q(0.05) q-nonpar (12); q-par (22) 4 Q(0.95) q-par (21) DB (4) 4 std chi-sq (14) DB (6) 8 corr fisher (6) 8 mean t-test (4); wilcoxon (6) DB (4) 8 median m-j (2); q-par (1); wilcoxon (3) DB (4) 8 Q(0.05) q-nonpar (12); q-par (22) DB (3) 8 Q(0.95) q-nonpar (12); q-par (21) 8 std chi-sq (8) DB (11) 16 mean t-test (8); wilcoxon (8) DB (13) 16 median wilcoxon (1) DB (7) 16 Q(0.05) q-nonpar (17); q-par (15) 16 Q(0.95) m-j (11); q-par (11) DB (5) 16 std chi-sq (2) DB (12) 32 mean t-test (3); wilcoxon (3) DB (14) 32 median m-j (1) DB (3) 32 Q(0.05) m-j (6); q-nonpar (17); q-par (5) 32 Q(0.95) m-j (18); q-par (6) 32 std chi-sq (2) DB (15) 64 mean DB (12) 64 median DB (1) 64 Q(0.05) m-j (6); q-nonpar (3) 64 Q(0.95) m-j (5) DB (11) 64 std chi-sq (1) DB (20) 128 mean DB (14) 128 Q(0.05) m-j (9) DB (1) 128 Q(0.95) m-j (5) DB (1) 128 std DB (28) 256 mean DB (10) 256 Q(0.05) DB (4) 256 Q(0.95) DB (10) 256 std DB (27)
Table 3.3 adds detail to the results from Section 3.2. For Pearson correlation DB does not produce CIs for $n=4$ and performs poorly for $n=8$. It performs poorly on the two extreme percentiles for $n\leq 32$ and in some cases on the other functionals, mostly for $n\leq 8$. For Q(0.05) and Q(0.95), we would do better by choosing q-par for the smaller sample sizes and m-j for sample size $n=32$. For the other functionals and $n\geq 8$ ($n\geq 16$ for Pearson correlation) there is no clear advantage of using baseline methods. Note that for Q(0.05) and Q(0.95) and $n=16,32$ B-n is as good as or better than baseline. However, for $n\leq 8$ even B-n is worse than the baseline, but less so than DB.
Table 3.3 adds detail to the results from Section 3.2. For Pearson correlation DB does not produce CIs for $n=4$ and performs poorly for $n=8$. It performs poorly on the two extreme percentiles for $n\leq 32$ and in some cases on the other functionals, mostly for $n\leq 8$. For Q(0.05) and Q(0.95), we would do better by choosing q-par for the smaller sample sizes and m-j for sample size $n=32$. For the other functionals and $n\geq 8$ ($n\geq 16$ for Pearson correlation) there is no clear advantage of using baseline methods. Table 3.3 shows the advantage of using B-n instead of DB for Q(0.05) and Q(0.95) for $n\leq 32$.
@@
Cases where B-n and baseline outperform each other. Results are for the two extreme percentiles and $n\leq 32$. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional. n functional baseline $\gg$ B-n B-n $\gg$ baseline 4 Q(0.05) q-nonpar (11); q-par (14) 4 Q(0.95) q-par (14) 8 Q(0.05) q-par (13) B-n (1) 8 Q(0.95) q-nonpar (4); q-par (13) B-n (4) 16 Q(0.05) q-nonpar (5); q-par (6) B-n (5) 16 Q(0.95) m-j (2); q-par (6) B-n (2) 32 Q(0.05) B-n (10) 32 Q(0.95) m-j (3); q-par (1) B-n (11)

### v3

@@ 3.1 When methods fail to produce a CI
There are only a few cases where a method fails to produce a CI. All bootstrap methods fail to produce a CI for Pearson correlation for $n=4$, due to division by zero variance. It can also happen for $n=8$, but rarely. BC and BCa do not produce a CI for the $5^{th}$ percentile for $n\leq 8$. For the $95^{th}$ percentile, B-t does not produce an interval in most cases when $n=4$ for the Laplace distribution. m-j is unable to produce CIs for small sample sizes and extreme percentiles ($5^{th}$ percentile for $n\leq 16$ and the $95^{th}$ percentile for $n\leq 8$). Method q-nonpar fails to produce the $95^{th}$ percentile for $n=4$ and $\alpha\leq 0.75$, for $n\in\{8,16\}$ and $\alpha\leq 0.25$, for $n=32$ and $\alpha\leq 0.05$ and for $n=64$ and $alpha=0.025$. When predicting CIs for the median, it fails at $n=4$ and $\alpha\leq 0.05$. And, although wilcoxon returns CIs for asymmetric distributions, they are not useful CIs for location. Also note that all cases for Bernoulli, where all $n$ had the same value, are removed, because most methods fail if there is no variability in the data.
@@ 3.2 Coverage of bootstrap methods
Table 3.2 shows a comparison of bootstrap methods in mean KL. As expected, coverage improves with sample size and the two extreme percentiles and standard deviation are the most difficult functionals. For Pearson correlation, mean, and median, DB is best. For the percentiles, B-n is best. And for standard deviation, B-t is best. B-n and B-t perform the best overall. Because coverage gets wors with smaller sample size, the overall results are biased towards methods that perform well on small $n$. DB and BCa, which are expected to perform best, perform relatively poorly for small $n$, but are best and second best for sample sizes $n\geq 64$.
@@
Mean KL coverage performance of bootstrap methods for one-sided CIs. The all column is across all combinations, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std B-n $\underline{0.078}$ $0.317$ $\underline{0.174}$ $\underline{0.087}$ $\underline{0.037}$ $0.021$ $0.014$ $0.009$ $0.006$ $0.026$ $0.004$ $\underline{0.024}$ $\underline{0.145}$ $0.208$ B-t $0.084$ $\underline{0.196}$ $0.269$ $0.107$ $0.049$ $0.023$ $0.019$ $0.014$ $0.008$ $\underline{0.013}$ $0.012$ $0.078$ $0.264$ $\underline{0.084}$ $B$ $0.112$ $0.249$ $0.189$ $0.218$ $0.101$ $0.046$ $0.033$ $0.022$ $0.017$ $0.031$ $0.045$ $0.159$ $0.234$ $0.145$ SB $0.118$ $0.452$ $0.267$ $0.152$ $0.060$ $0.027$ $0.017$ $0.011$ $0.002$ $0.022$ $0.003$ $0.064$ $0.230$ $0.308$ DB $0.134$ $0.536$ $0.340$ $0.193$ $0.057$ $\underline{0.008}$ $\underline{0.005}$ $\underline{0.003}$ $\underline{0.000}$ $\underline{0.013}$ $\underline{0.002}$ $0.125$ $0.401$ $0.188$ PB $0.157$ $0.610$ $0.364$ $0.219$ $0.073$ $0.028$ $0.018$ $0.011$ $0.002$ $0.025$ $0.003$ $0.109$ $0.377$ $0.330$ BC $0.161$ $0.628$ $0.359$ $0.228$ $0.092$ $0.022$ $0.013$ $0.008$ $0.001$ $0.049$ $0.017$ $0.169$ $0.387$ $0.246$ BCa $0.161$ $0.632$ $0.359$ $0.236$ $0.092$ $0.018$ $0.010$ $0.005$ $0.001$ $0.058$ $0.017$ $0.201$ $0.379$ $0.218$
@@
Threshold-based coverage performance of bootstrapping on one-sided CIs. The table shows the percentage of experiments where a method does not meet the liberal criterion $25\times\text{KL}(0.945,0.95)$. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std DB $\underline{0.30}$ $0.67$ $\underline{0.49}$ $\underline{0.46}$ $0.44$ $\underline{0.12}$ $\underline{0.07}$ $\underline{0.03}$ $\underline{0.00}$ $\underline{0.18}$ $\underline{0.05}$ $0.46$ $0.50$ $\underline{0.42}$ B-n $0.40$ $0.78$ $0.65$ $0.56$ $\underline{0.36}$ $0.28$ $0.20$ $0.13$ $0.20$ $0.34$ $0.11$ $\underline{0.28}$ $0.46$ $0.80$ BCa $0.41$ $0.89$ $0.70$ $0.62$ $0.41$ $0.23$ $0.21$ $0.06$ $\underline{0.00}$ $0.32$ $0.33$ $0.68$ $\underline{0.41}$ $0.47$ SB $0.45$ $0.83$ $0.63$ $0.57$ $0.49$ $0.40$ $0.28$ $0.14$ $0.07$ $0.28$ $0.12$ $0.49$ $0.57$ $0.94$ BC $0.46$ $0.89$ $0.72$ $0.68$ $0.50$ $0.34$ $0.24$ $0.11$ $\underline{0.00}$ $0.31$ $0.33$ $0.57$ $0.73$ $0.54$ PB $0.47$ $0.82$ $0.66$ $0.63$ $0.51$ $0.39$ $0.28$ $0.16$ $0.07$ $0.32$ $0.10$ $0.49$ $0.59$ $0.94$ B-t $0.47$ $\underline{0.54}$ $0.62$ $0.55$ $0.51$ $0.43$ $0.41$ $0.32$ $0.23$ $0.30$ $0.35$ $0.59$ $0.67$ $0.57$ $B$ $0.57$ $0.80$ $0.75$ $0.64$ $0.60$ $0.50$ $0.46$ $0.41$ $0.27$ $0.37$ $0.59$ $0.76$ $0.81$ $0.50$
@@
Threshold-based coverage performance of bootstrapping on two-sided CIs. The table shows the percentage of experiments where a method does not meet the liberal criterion $25\times\text{KL}(0.945,0.95)$. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std DB $\underline{0.33}$ $0.82$ $\underline{0.53}$ $\underline{0.58}$ $\underline{0.53}$ $\underline{0.11}$ $\underline{0.03}$ $\underline{0.03}$ $\underline{0.00}$ $0.29$ $\underline{0.07}$ $0.43$ $\underline{0.40}$ $\underline{0.57}$ SB $0.44$ $1.00$ $0.68$ $0.71$ $0.62$ $0.23$ $0.12$ $0.08$ $0.10$ $0.44$ $0.14$ $0.43$ $0.42$ $0.82$ B-n $0.46$ $0.97$ $0.82$ $0.76$ $0.56$ $0.29$ $0.12$ $0.08$ $0.40$ $0.46$ $0.23$ $\underline{0.35}$ $0.43$ $0.80$ BCa $0.47$ $1.00$ $0.89$ $0.68$ $0.59$ $0.33$ $0.11$ $0.06$ $\underline{0.00}$ $0.42$ $0.25$ $0.62$ $\underline{0.40}$ $0.74$ PB $0.47$ $1.00$ $0.76$ $0.76$ $0.62$ $0.30$ $0.12$ $0.08$ $0.20$ $0.44$ $0.19$ $0.47$ $0.47$ $0.82$ BC $0.48$ $1.00$ $0.92$ $0.73$ $0.58$ $0.36$ $0.09$ $0.06$ $\underline{0.00}$ $0.43$ $0.25$ $0.57$ $0.45$ $0.77$ B-t $0.66$ $\underline{0.53}$ $0.74$ $0.80$ $0.74$ $0.62$ $0.58$ $0.56$ $0.40$ $\underline{0.21}$ $0.76$ $1.00$ $1.00$ $0.65$ $B$ $0.80$ $1.00$ $1.00$ $0.92$ $0.83$ $0.76$ $0.64$ $0.59$ $0.60$ $0.47$ $0.95$ $1.00$ $1.00$ $0.79$
@@
Four of the six confidence levels in the experiments can be used to derive results for the two most common two-sided CIs (95%, 90%). Table 3.2 shows how often the methods fail to meet the liberal criterion for these two-sided intervals. These and results for other thresholds and KL are similar for one-sided and two-sided intervals. Most bootstrap methods meet the liberal criterion in almost all cases when $n\geq 128$ ($n\geq 64$ for DB), but even at $n=256$ there are still two cases where even DB does not (standard deviation for log-normal). Excluding these, DB meets the very liberal criterion for all experiments for $n\geq 64$.
@@ 3.3 Coverage comparison with baseline methods
All of the results in this section are for one-sided CIs. Results for two-sided CIs are similar. Starting with the premise that when limited to a single method, DB is the best choice, this section takes a closer look at where DB is clearly outperformed by another method. The criterion used is that method A outperforms method B if B does not meet the liberal criterion and A is at least an order of magnitude better than method B. If a method is not able to produce CIs, it is outperformed by any method that can.
@@
Cases where DB and baseline outperform each other. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6, and the mean, where there are 48. n functional baseline $\gg$ DB DB $\gg$ baseline 4 corr fisher (6) 4 mean a-c (2); c-p (1); t-test (11); wilcoxon (1) DB (2) 4 median m-j (5); q-par (3) 4 Q(0.05) q-nonpar (12); q-par (22) 4 Q(0.95) q-par (21) DB (4) 4 std chi-sq (14) DB (5) 8 corr fisher (6) 8 mean a-c (7); c-p (5); t-test (10); wilcoxon (3) DB (3) 8 Q(0.05) q-nonpar (12); q-par (22) DB (3) 8 Q(0.95) q-nonpar (12); q-par (21) 8 std chi-sq (7) DB (9) 16 mean a-c (4); c-p (3); t-test (4) DB (9) 16 Q(0.05) q-nonpar (17); q-par (15) 16 Q(0.95) m-j (10); q-par (11) DB (4) 16 std chi-sq (2) DB (9) 32 mean a-c (1); c-p (1); t-test (3) DB (8) 32 Q(0.05) m-j (6); q-nonpar (17); q-par (5) 32 Q(0.95) m-j (17); q-par (6) 32 std DB (12) 64 mean a-c (1); c-p (1) DB (7) 64 Q(0.05) m-j (1); q-nonpar (3) 64 Q(0.95) DB (5) 64 std chi-sq (1) DB (18) 128 mean a-c (2); c-p (1); t-test (1) DB (6) 128 Q(0.05) m-j (2) DB (1) 128 std DB (24) 256 mean DB (4) 256 std DB (27)
Table 3.3 adds detail to the results from Section 3.2. For Pearson correlation DB does not produce CIs for $n=4$ and performs poorly for $n=8$. It performs poorly on the two extreme percentiles for $n\leq 32$ and in some cases on the other functionals, mostly for $n\leq 8$. For Q(0.05) and Q(0.95), a better choice is q-par for the smaller sample sizes and m-j for sample size $n=32$. For the other functionals and $n\geq 8$ ($n\geq 16$ for Pearson correlation) there is no clear advantage of using baseline methods. Note that for Q(0.05) and Q(0.95) and $n=16,32$ B-n is as good as or better than baseline. Table 3.3 shows the advantage of using B-n instead of DB for Q(0.05) and Q(0.95) for $n\leq 32$.
@@
Cases where B-n and baseline outperform each other. Results are for the two extreme percentiles and $n\leq 32$. The number in the parentheses is the number of combinations where a method outperforms. There are a total of 36 combinations for every pair of $n$ and functional. n functional baseline $\gg$ B-n B-n $\gg$ baseline 4 Q(0.05) q-nonpar (11); q-par (14) 4 Q(0.95) q-par (14) 8 Q(0.05) q-par (11) B-n (1) 8 Q(0.95) q-nonpar (4); q-par (10) B-n (3) 16 Q(0.05) q-nonpar (5); q-par (6) B-n (5) 16 Q(0.95) m-j (1); q-par (6) 32 Q(0.05) B-n (4) 32 Q(0.95) m-j (2); q-par (1) B-n (6)